# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ 3. $(\mathrm{3})$
We will first prove $(1)\Leftrightarrow(2)$ and then $(2)\Leftrightarrow(3)$.
(1$\to$(2): Can be obtained just by appropriate substitution in Lemma C.6.
(2$\to$(1): Using Lemma C.6, we get $\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})\Rightarrow\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi_{\nu^{1/2}\mathrm{P}\nu^{1/2}}).$ Now using Lemma C.5, we get $\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi_{\nu^{1/2}\mathrm{P}\nu^{1/2}})\Rightarrow\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi).$
(2$\to$(3): Using Lemma C.4, we know that $\Pi\Pi_{\nu}=\Pi\Rightarrow\mathrm{P}=\mathrm{P}\Pi_{\nu}$. Hence
@@
Using Lemma C.1, $(\Pi_{\nu}-\mathrm{P})\zeta(\Pi_{\nu}-\mathrm{P})=0\Rightarrow(\Pi_{\nu}-\mathrm{P})\zeta=\zeta(\Pi_{\nu}-\mathrm{P})=0$ and so
$\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})$ $\Rightarrow(\mathrm{I}-\Pi_{\nu}+\mathrm{P})\zeta(\mathrm{I}-\Pi_{\nu}+\mathrm{P})=\zeta\Rightarrow\zeta\in\mathcal{S}(\mathrm{I}-\Pi_{\nu}+\mathrm{P}).$
(3$\to$(2): Using Lemma C.4, we know that $\Pi\Pi_{\nu}=\Pi\Rightarrow\mathrm{P}=\mathrm{P}\Pi_{\nu}$. Hence
$\zeta\in\mathcal{S}(\mathrm{I}-\Pi_{\nu}+\mathrm{P})$ $\Rightarrow\zeta=(\mathrm{I}-\Pi_{\nu}+\mathrm{P})\zeta$
@@
$\Rightarrow\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{S}(\mathrm{P}).$
which completes the proof. ∎
Theorem C.2.
@@ 1. $(\mathrm{1})$
    $\zeta\in\mathcal{P}(\mathrm{P})\Rightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$ and so $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$.
@@ 2. $(\mathrm{2})$
We will prove $\mathrm{(1)}$ and then $\mathrm{(2)}$.
[…]
@@
[…]
Appendix D Minimizing the max norm
Lemma D.1.
For any $\nu\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$, $\|\Pi\nu\Pi\|_{\infty}\leq\|\nu\|_{\infty}$.
Proof.
Note that $\langle\phi|\Pi\nu\Pi|\phi\rangle=\langle\phi|\Pi|\phi\rangle\frac{\langle\phi|\Pi}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\nu\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\leq\frac{\langle\phi|\Pi}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\nu\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}.$ Beginning with the definition of max-norm and using this inequality, we get
$\|\Pi\nu\Pi\|_{\infty}$ $=\max_{|\phi\rangle}\langle\phi|\Pi\nu\Pi|\phi\rangle\leq\max_{|\phi\rangle}\frac{\langle\phi|\Pi}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\nu\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\max_{|\phi\rangle:|\phi\rangle=\Pi|\phi\rangle}\langle\phi|\nu|\phi\rangle\stackrel{{\scriptstyle(b)}}{{\leq}}\max_{|\phi\rangle}\langle\phi|\nu|\phi\rangle\stackrel{{\scriptstyle(c)}}{{=}}\|\nu\|_{\infty}.$
Here, $(a)$ follows from the fact that $\left\{\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}:|\phi\rangle\right\}=\left\{\phi:\Pi|\phi\rangle=|\phi\rangle\right\}$. $(b)$ is obtained because the restriction $\{|\phi\rangle:|\phi\rangle=\Pi|\phi\rangle\}$ is removed and $(c)$ is obtained from the definition of max-norm. ∎
Lemma D.2.
For any $\sigma\in\mathcal{S}(\mathcal{H})$ and projector $\mathrm{P}$ such that $\mathrm{P}\Pi_{\sigma}=\mathrm{P}$ then
$\min_{\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\frac{1}{\mathrm{Tr}\left(\mathrm{P}\sigma\right)}.$
@@ Appendix D Minimizing the max norm
We will prove this in 3 steps by showing the following as equal.
@@
Now, note that $\mathrm{Tr}\left(\psi\sigma\right)$ is constant and so using Lemma D.1, we get $\left\|\frac{\Pi_{\sigma}\psi\Pi_{\sigma}}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}$, so
@@
Substituting $\psi^{\prime}=\Pi_{\sigma}\psi\Pi_{\sigma}$, and from Theorem C.2 in Appendix C, we know that $\psi^{\prime}\in\{\Pi_{\sigma}\psi\Pi_{\sigma}:\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}\subseteq\mathcal{P}(\mathrm{P})$. Hence,
$\min_{\psi^{\prime}\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi^{\prime}}{\mathrm{Tr}\left(\psi^{\prime}\sigma\right)}\right\|_{\infty}\leq\min_{\psi^{\prime}\in\{\Pi_{\sigma}\psi\Pi_{\sigma}:\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}}\left\|\frac{\psi^{\prime}}{\mathrm{Tr}\left(\psi^{\prime}\sigma\right)}\right\|_{\infty}=\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\Pi_{\sigma}\psi\Pi_{\sigma}}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (72)
Hence
$\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (73)
But, from Theorem C.2 in Appendix C, we know that $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$, hence
@@
2$\to$3: Taking $\phi=\frac{\psi}{\|\psi\|_{\infty}}$, we get the constrain as $\phi\in\mathcal{P}(\mathrm{P})$ and $\|\phi\|_{\infty}=1$ and minimum as
@@
Lemma D.3.
For any $\rho,\sigma$ such that $\Pi_{\rho}=\Pi_{\sigma}$, $10<c_{r}<$ 1,
@@
Here, for any $\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0$, we have
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)\stackrel{{\scriptstyle\Delta}}{{=}}\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2})\}.$
More detailed studey of the function is given in given in Appendix E.
@@
=$=\left\{\max C:C\left(c_{r}\phi_{\max}+(1-c_{r})\phi_{\min}\right)\leq\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2}\ \text{for some }{\phi_{\max}\in\mathcal{S}(\mathrm{T}^{\max}),\phi_{\min}\in\mathcal{S}(\mathrm{T}^{\min})}\right\}^{-1}$
@@
thus completing the proof. ∎
Appendix E Definition and properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$
@@
It denotes the largest possible trace of any matrix $\psi\in\mathcal{P}(\Pi_{1}+\Pi_{2})$ with property $\psi\leq\sigma$, $\mathrm{Tr}\left(\psi\Pi_{1}\right)=r\mathrm{Tr}\left(\psi\right)$ and $\Pi_{1}\psi\Pi_{2}=0$.
We will take cases when $\Pi_{1}+\Pi_{2}\leq\Pi_{\sigma}$. If this not the cases, $\Pi_{1}$ (and $\Pi_{2}$) can just be substituted by the projector corresponding to the subspace spanned by the intersection subspace of $\Pi_{1}$ (and $\Pi_{2}$) and $\Pi_{\sigma}$.
@@
  • $\Upsilon_{\Pi_{1},\Pi_{2}}(k\sigma,r)=k\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)$
@@
Some closed-form expressions: These closed form expression obtained for specific cases help in obtaining closed form expression in $\mathcal{C}3$ in Theorem III.4.
@@
  • If $\Pi_{\sigma}>\Pi_{1}+\Pi_{2}$, the substitution $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma^{1/2}\Pi_{\sigma^{-1/2}(\Pi_{1}+\Pi_{2})\sigma^{-1/2}}\sigma^{1/2},r\right)$ simplifies it as $\sigma^{1/2}\Pi_{\sigma^{-1/2}(\Pi_{1}+\Pi_{2})\sigma^{-1/2}}\sigma^{1/2}\in\mathcal{P}(\Pi_{1}+\Pi_{2})$. From here on-words, we can proceed with the assumption that $\Pi_{1}+\Pi_{2}=\Pi_{\sigma}$.
Lemma E.1.
@@
$\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\begin{cases}\displaystyle\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r},&\displaystyle r\leq\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{\mathrm{Tr}\left(\sigma\right)},\vspace{1mm}\\ \displaystyle\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},&\text{else}.\end{cases}.$
@@
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)=\{\max c:c(r\Pi_{1}+(1-r)\Pi_{2})\leq\sigma\}.$
@@
Hence $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}$. ∎
Appendix F Proof of Theorem IV.1
We will use frequently here that $\mathrm{Tr}\left(\Gamma\nu\right)=0\Leftrightarrow\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\nu})$ (See Lemma C.2 in Appendix C for proof). Starting with the definition of set $\mathcal{M}$ and the condition $e(\Lambda)=0$, the goal to find set of all ${\Lambda}$ such that $\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ and $e(\Lambda)=0$. So,
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\{\Lambda:e(\Lambda)=0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$
=$=\left\{\Lambda:\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}=0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (76)
The previous step follows from using the definition of $e(\Lambda)$. Now, for the fraction to be 00, numerator must be zero and denominator must remain non-zero. Hence, we have

### v2

@@ 3. $(\mathrm{3})$
We will first prove $(1)\Leftrightarrow(2)$, and then $(2)\Leftrightarrow(3)$.
[…]
@@
Using Lemma B.1, $(\Pi_{\nu}-\mathrm{P})\zeta(\Pi_{\nu}-\mathrm{P})=0\Rightarrow(\Pi_{\nu}-\mathrm{P})\zeta=\zeta(\Pi_{\nu}-\mathrm{P})=0$ and so
$\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})$ $\Rightarrow(\mathrm{I}-\Pi_{\nu}+\mathrm{P})\zeta(\mathrm{I}-\Pi_{\nu}+\mathrm{P})=\zeta\Rightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}+\mathrm{P}).$
[…]
$\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}+\mathrm{P})$ $\Rightarrow\zeta=(\mathrm{I}-\Pi_{\nu}+\mathrm{P})\zeta$
@@
$\Rightarrow\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P}).$
Thus proved. ∎
Theorem B.2.
@@ 1. $(\mathrm{1})$
    $\zeta\in\mathcal{P}(\mathrm{P})\Rightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$ and so $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$, and
@@ 2. $(\mathrm{2})$
We will prove $\mathrm{(1)}$, and then $\mathrm{(2)}$.
(1): Beginning with the definition of $\mathcal{P}(\mathrm{P})$, we get $\zeta\in\mathcal{P}(\mathrm{P})\Rightarrow\zeta=\mathrm{P}\zeta$. Now, multiplying both sides by $\Pi_{\sigma}$ and simplifying further, we obtain $\Pi_{\sigma}\zeta=\Pi_{\sigma}\mathrm{P}\zeta=\mathrm{P}\zeta$. So, we get $(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\zeta=\zeta$ and thus $\zeta=\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$. Now, note that $\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\ \forall\ \zeta\in\mathcal{P}(\mathrm{P})$, so $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$.
@@
Similarly, it can be shown that $\zeta\mathrm{P}=\zeta\Pi_{\sigma}\mathrm{P}$. So, we get $\Pi_{\sigma}\zeta\Pi_{\sigma}=\mathrm{P}\Pi_{\sigma}\zeta\Pi_{\sigma}\mathrm{P}$. Now, by the definition of $\mathcal{P}(\mathrm{P})$, we get the stated result as $\Pi_{\sigma}\zeta\Pi_{\sigma}\in\mathcal{P}(\mathrm{P})$. Further, this is true for all $\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$, so we get $\{\Pi_{\sigma}\zeta\Pi_{\sigma}:\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}\subseteq\mathcal{P}(\mathrm{P})$. ∎
Appendix C Minimizing the max norm
Lemma C.1.
For any $\sigma\in\mathcal{S}(\mathcal{H})$, and projector $\mathrm{P}$ such that $\mathrm{P}\Pi_{\sigma}=\mathrm{P}$, then
$\min_{\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\frac{1}{\mathrm{Tr}\left(\mathrm{P}\sigma\right)},$
@@ Appendix C Minimizing the max norm
We will prove the statement in lemma in 3 steps by showing the following as equal.
@@
Now, using $\left\|{\Pi_{\sigma}\psi\Pi_{\sigma}}\right\|_{\infty}\leq\left\|{\psi}\right\|_{\infty}$, we get
@@
Substituting $\psi^{\prime}=\Pi_{\sigma}\psi\Pi_{\sigma}$, and from Theorem B.2 in Appendix B, we know that $\psi^{\prime}\in\{\Pi_{\sigma}\psi\Pi_{\sigma}:\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}\subseteq\mathcal{P}(\mathrm{P})$. Hence,
$\min_{\psi^{\prime}\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi^{\prime}}{\mathrm{Tr}\left(\psi^{\prime}\sigma\right)}\right\|_{\infty}\leq\min_{\psi^{\prime}\in\{\Pi_{\sigma}\psi\Pi_{\sigma}:\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}}\left\|\frac{\psi^{\prime}}{\mathrm{Tr}\left(\psi^{\prime}\sigma\right)}\right\|_{\infty}=\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\Pi_{\sigma}\psi\Pi_{\sigma}}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}$ (50)
$\Rightarrow\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (51)
But, from Theorem B.2 in Appendix B, we know that $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$, hence
@@
[…]
@@
Lemma C.2.
For any $\rho,\sigma$ such that $\Pi_{\rho}=\Pi_{\sigma}$, $0<c_{r}<1$,
@@
Here, $\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)$ is defined as
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)\stackrel{{\scriptstyle\Delta}}{{=}}\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2}),\text{ with }\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0\}.$
More detailed properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$ are given in Appendix D.
@@
$=\left\{\max C:C\left(c_{r}\phi_{\max}+(1-c_{r})\phi_{\min}\right)\leq\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2}\ \text{for some }{\phi_{\max}\in\mathcal{S}(\mathrm{T}^{\max}),\phi_{\min}\in\mathcal{S}(\mathrm{T}^{\min})}\right\}^{-1}$
@@
where last step follows from definition of $\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}$ in Appendix D. Thus, completing the proof. ∎
Appendix D Definition and properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$
@@
It denotes the largest value of trace of matrix $\psi\in\mathcal{P}(\Pi_{1}+\Pi_{2})$, having the properties $\psi\leq\sigma$, $\mathrm{Tr}\left(\psi\Pi_{1}\right)=r\mathrm{Tr}\left(\psi\right)$ and $\Pi_{1}\psi\Pi_{2}=0$.
In the rest of discussion, we consider the case when $\Pi_{1}+\Pi_{2}\leq\Pi_{\sigma}$. If this not the cases, $\Pi_{1}$ (and $\Pi_{2}$) can just be substituted by the projector corresponding to the subspace spanned by the intersection subspace of $\Pi_{1}$ (and $\Pi_{2}$) and $\Pi_{\sigma}$. The following two properties can be obtained by substituting the definition of $\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)$.
@@
  • $\Upsilon_{\Pi_{1},\Pi_{2}}(k\sigma,r)=k\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)$,
@@
Some closed-form expressions: These closed form expression obtained for specific cases help in obtaining closed form expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$ in case $\mathcal{C}3$ in Theorem 4.
@@
  • If $\Pi_{\sigma}>\Pi_{1}+\Pi_{2}$, substituting $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma^{1/2}\Pi_{\sigma^{-1/2}(\Pi_{1}+\Pi_{2})\sigma^{-1/2}}\sigma^{1/2},r\right)$ simplifies it as $\sigma^{1/2}\Pi_{\sigma^{-1/2}(\Pi_{1}+\Pi_{2})\sigma^{-1/2}}\sigma^{1/2}\in\mathcal{P}(\Pi_{1}+\Pi_{2})$.
From here on-words, we can proceed with the assumption that $\Pi_{1}+\Pi_{2}=\Pi_{\sigma}$.
Lemma D.1.
@@
$\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\begin{cases}\displaystyle\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r},&\displaystyle r\leq\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{\mathrm{Tr}\left(\sigma\right)},\vskip 2.84526pt\\ \displaystyle\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},&\text{otherwise}.\end{cases}.$
@@
From the definition, $\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)=\{\max c:c(r\Pi_{1}+(1-r)\Pi_{2})\leq\sigma\}.$ Hence,
@@
So, we obtain $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}$. ∎
Appendix E Proof of Theorem 5
We will use frequently here that $\mathrm{Tr}\left(\Gamma\nu\right)=0\Leftrightarrow\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\nu})$ (See Lemma B.1 in Appendix B for proof). Starting with the definition of set $\mathcal{M}$ and the condition $e(\Lambda)=0$, the goal to find set of all $\Lambda$} such that $\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ and $e(\Lambda)=0$. So,
$\mathcal{E}_{s}(\rho,\sigma,p)$ $=\{\Lambda:e(\Lambda)=0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$
$=\left\{\Lambda:\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}=0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (54)
The previous step follows from using the definition of $e(\Lambda)$. Now, for a fraction to be 0, numerator must be zero and denominator must remain non-zero. Hence, we get